i need help solving a chapter problem.
its a word problem that many students have trouble with.
So far i have gotten most of it until it said to find a number to make to solution exist.
this is what i have.
x+y=45
11x+10y= ?
i have tried 100, 80, & 81 but none of them worked , it was all no solution.
i know how to graph the problems & solve them alebraically . Its just finding a number to make it a solution is what i am having trouble with .
please help D:
if you need more info about the problem please email me.

Answer by rwm(914)

You can't find numbers. You now have three unknowns with only two equations.
x+y=45
y=45-x
11x+10y= ?
11x+10*(45-x)=
11x+450-10x=x+450
11x+10y=x+450
10x+10y=450
x+y=45
You can only solve in terms of x or in terms of y.